Q1)Determine the [H3O+] in a 0.256 M HClO solution. The Ka of HClO is 2.8  10–9.

Q2)Find the percent ionization of a 0.373 M HF solution. The Ka for HF is 3.4  10–5.

thanks!
Answer #1

Не ощ) = clo (aq) + H+ (aq) guitial: 0.256 final: (01256x) - Lets assume x to be small x ka is small. as 0.256-x ~ 0.250 KeG H E (aq) = hit coul + F Caps 0.373 (0.373-x) x x I will be smell Lets assume a as be ka is small small. ... 0.373-x = 0.37